OpenStudy (anonymous):

What is the product of \[5\sqrt{25d} \dot\ 2d\sqrt{3}\]

OpenStudy (anonymous):

\[5\sqrt{25d} \dot\ 2d\sqrt{3}\]

OpenStudy (anonymous):

@Loujoelou

OpenStudy (anonymous):

Well, first let's see if we can take anything out of the square roots. Do u know how to take numbers and variables out of square roots? Do u know how to simplify roots?

OpenStudy (anonymous):

no clue

OpenStudy (anonymous):

is it \[50d\] \sqrt{3d}

OpenStudy (anonymous):

50d sqrt(3d)

OpenStudy (anonymous):

K :p